OpenStudy (anonymous):

Estimate and round to the nearest whole number 6.07x3.29=???

OpenStudy (anonymous):

better use calculator

OpenStudy (saifoo.khan):

i used.

OpenStudy (saifoo.khan):

it said, 19.97 so 20.

OpenStudy (anonymous):

Thanks :)

OpenStudy (saifoo.khan):

welks

OpenStudy (anonymous):

So I dont need to write 19.97?

OpenStudy (saifoo.khan):

NO.

OpenStudy (saifoo.khan):

whole number doesnt contain any decimals.

myininaya (myininaya):

ok keep in mind we need to move the decimal over 4 times so we have\[607 \cdot 329=7(329)+0(3290)+6(32900)=2303+0+197400=199703\] \[6.07 \cdot 3.29=19.9703\] nearest whole number is 20

OpenStudy (anonymous):

So I can just solve the problem and round it? Thanks btw

OpenStudy (saifoo.khan):

yesss!